OpenStudy (zaxoanl):

a trough is filled with water weighing 62.4 pounds per cubic foot and its vertical ends have the shape of the parabolic region. find the hydrostatic force on one end of the trough.

OpenStudy (zaxoanl):

so F=PA P=distance*density \[(x-h)^2+(y-k)^2=r^2\] \[x=\sqrt{16-(y-4)^2}\] \[62.4\int\limits_{0}^{4}2\sqrt{16-(y-4)^2(4-y)}dy\]

OpenStudy (zaxoanl):

not sure if i set it up correctly and how do you integrate it

OpenStudy (irishboy123):

you're using the equation of a circle,but isn't it parabolic? maybe you have a link or something to share?